Find the variation constant, and write a formula that expresses the indicated variation.
c varies inversely as d, and c  5 when d  2.
I'm not grasping the concept here, please help.

Answer by gonzo(654)

* In inverse variation, when one variable increases the other decreases in proportion so that the product remains the same always.
-----
your problem states that c varies inversely as d.
not sure of the symbols, but i understand that part to mean that c = 5 when d = 2.
-----
formula for constant of inverse variation is xy = k or x = k/y, or y = k/x
k is the constant of variation.
-----
in your problem, that formula would appear as cd = k, or c = k/d or d = k/c.
if c is 5, and d is 2, the formula becomes (5*2) = k = 10
-----
constant variation appears to be 10.
c = k/d becomes 5 = 10/2 = 5.
d = k/c becomes 2 = 10/5 = 2.
-----
it works out if i understood your problem correctly.
k = 10.
-----
the formula that expresses this relationship would be cd = 10.
that would equate to either:
c = 10/d
or
d = 10/c
-----